OpenStudy (anonymous):

-WILL MEDAL- Write the following in standard form: (0, -4), (-1, 5). Please double check my answer in the comments. Thanks

OpenStudy (anonymous):

Slope = (5+4)/(0+1) = 9/1 = 9 y-Intercept = y = 9x + b -4 = 9(0) + b => -4 = b

OpenStudy (anonymous):

y = 9x - 4 -9x + y = -4

OpenStudy (kainui):

Are you sure the slope is (5+4)/(0+1) ?

OpenStudy (anonymous):

Oh I see, -1?

OpenStudy (anonymous):

So 9x + y = -4?

OpenStudy (kainui):

One way is to check yourself, plug in both points to your line, I'll do the first one: 9*(0)+(-4)=-4 That checks out, now try the other point.

OpenStudy (anonymous):

I did the math on paper and that's what I got... dang signs -.-

OpenStudy (kainui):

Hey, at least both points work out, must mean it's correct. =P